A population consists of five numbers 1, 2, 3, 4, 5 and 6. Suppose samples of size 2 are drawn from this



population.



(a) Find the mean and variance of the population



(b) Describe the sampling distribution of the sample means.



(c) Find the mean and variance of the sampling distribution of the sample means.

1
Expert's answer
2022-05-18T06:55:49-0400

(a) We have population values 1,2,3,4,5,6, population size N=6 and sample size n=2.

Mean of population "(\mu)" = "\dfrac{1+2+3+4+5+6}{6}=3.5"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{1}{6}(6.25+2.25+0.25"




"+0.25+2.25+6.25)=\dfrac{17.5}{6}"


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{17.5}{6}}\approx1.707825"


(b) The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_2=15."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 1,2 & 3/2 \\
 \hdashline
 2 & 1,3 & 4/2 \\
 \hdashline
 3 & 1,4 & 5/2\\
 \hdashline
 4 & 1,5 & 6/2 \\
 \hdashline
 5 & 1,6 & 7/2 \\
 \hdashline
 6 & 2,3 & 5/2 \\
 \hdashline
 7 & 2,4 & 6/2 \\
 \hdashline
 8 & 2,5 & 7/2 \\
 \hdashline
 9 & 2,6 & 8/2 \\
 \hdashline
 10 & 3,4 & 7/2 \\
 \hdashline
 11 & 3,5 & 8/2 \\
 \hdashline
 12 & 3,6 & 9/2 \\
 \hdashline
 13 & 4,5 & 9/2 \\
 \hdashline
 14 & 4,6 & 10/2 \\
 \hdashline
 15 & 5,6 & 11/2 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X})& \bar{X}^2f(\bar{X}) 
\\ \hline
3/2 & 1/15 & 3/30 & 9/60 \\
 \hdashline
4/2 & 1/15 & 4/30 & 16/60 \\
 \hdashline
5/2 & 2/15 & 10/30 & 50/60 \\
 \hdashline
6/2 & 2/15 & 12/30 & 72/60 \\
 \hdashline
7/2 & 3/15 & 21/30 & 147/60 \\
 \hdashline
8/2 & 2/15 & 16/30 & 128/60 \\
 \hdashline
9/2 & 2/15 & 18/30 & 162/60 \\
 \hdashline
10/2 & 1/15 & 10/30 & 100/60 \\
 \hdashline
11/2 & 1/15 & 11/30 & 121/60 \\
 \hdashline
\end{array}"



(c) Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{105}{30}=3.5=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{805}{60}-(3.5)^2=\dfrac{7}{6}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{7}{6}}\approx1.080123"